OpenStudy (ny,ny):

how do you find x-intercept? in one question the y-intercept was (0,30) and the x-intercept was (-30,0) does that mean the x-intercept is always the opposite of the y-intercept?

OpenStudy (anonymous):

you mean if you have straight line and you are trying to find it's interceptions with x-axis and y-axis? NO THEY ARE NOT OPPOSITE line equation: y=k*x+q x-axis interception = -q/k y-axis interception = q

OpenStudy (precal):

no when x=0 what is y? this is how you find the y intercept and when y=0 what is x? this is how you find the x intercept

OpenStudy (ny,ny):

thank you!!

OpenStudy (precal):

anytime :)